### 13 Mar 2026 v2

# Beyond Convolution: A Taxonomy of Structured Operators for Learning-Based Image Processing

Abstract
The convolution operator is the fundamental building block of modern convolutional neural networks (CNNs), owing to its simplicity, translational equivariance, and efficient implementation. However, its structure as a fixed, linear, locally-averaging operator limits its ability to capture structured signal properties such as low-rank decompositions, adaptive basis representations, and non-uniform spatial dependencies. This paper presents a systematic taxonomy of operators that extend or replace the standard convolution in learning-based image processing pipelines. We organise the landscape of alternative operators into five families: (i) decomposition-based operators, which separate structural and noise components through singular value or tensor decompositions; (ii) adaptive weighted operators, which modulate kernel contributions as a function of spatial position or signal content; (iii) basis-adaptive operators, which optimise the analysis bases together with the network weights; (iv) integral and kernel operators, which generalise the convolution to position-dependent and non-linear kernels; and (v) attention-based operators, which relax the locality assumption entirely. For each family, we provide a formal definition, a discussion of its structural properties with respect to the convolution, and a critical analysis of the tasks for which the operator is most appropriate. We further provide a comparative analysis of all families across relevant dimensions—linearity, locality, equivariance, computational cost, and suitability for image-to-image and image-to-label tasks—and outline the open challenges and future directions of this research area.
Beyond Convolution: A Taxonomy of Structured Operators for Learning-Based Image Processing
Simone Cammarasana
CNR-IMATI, Via De Marini 6, Genova, Italy
simone.cammarasana@ge.imati.cnr.it
Keywords: Convolutional neural networks, Alternative operators, Singular value decomposition, Weighted convolution, F-transform, Attention mechanism, Image processing, Taxonomy.
1 Introduction
Deep learning has established itself as the dominant paradigm for signal and image processing, achieving state-of-the-art performance on tasks ranging from image restoration [ZZC+17, DFK+06] and super-resolution [DLH+16] to semantic segmentation and medical image analysis [CP26]. At the heart of nearly every successful architecture lies the convolution operator, which extracts local features by sliding a fixed, learned kernel over the input signal. Despite its success, the convolution operator is fundamentally a weighted local average: it applies a fixed linear combination to each neighbourhood of pixels, sharing the same weights across all spatial positions.
This structural simplicity, while computationally advantageous, has well-known limitations. First, the convolution treats all spatial positions uniformly, making it insensitive to local signal structure such as edges, textures, or noise patterns. Second, its linear nature prevents it from performing operations that are inherently structural, such as separating a signal into low-rank and high-rank components, or projecting onto an application-specific basis. Third, the fixed kernel size imposes a rigid locality prior that is inappropriate for tasks requiring global context or multi-scale reasoning.
A growing body of work addresses these limitations by proposing operators that enrich or replace the convolution while remaining compatible with the back-propagation training framework. These contributions are distributed across multiple communities—signal processing, numerical linear algebra, fuzzy mathematics, and deep learning—making it difficult to obtain a unified view of the design space. Our work fills this gap by providing a systematic taxonomy of structured operators as alternatives or enhancements to convolution in learning-based image processing.
Motivating examples
The limitations of standard convolution are well illustrated by concrete tasks. In image denoising, applying the singular value decomposition (SVD) to local patches explicitly separates the structured (low-rank) signal component from the unstructured noise component [GZZ+14, CP22, HLL18]. In image super-resolution, interpolating bases such as B-splines encode smoothness priors that a fixed kernel cannot capture efficiently [DLH+16, SCH+16]. In classification, optimising the density function applied to the convolution kernel yields a spatially non-uniform weighting that improves convergence and accuracy [CP25b, CP25c]. In fuzzy signal processing, the F-transform with adaptive membership functions defines a projection operator whose bases are data-driven rather than fixed [CP24, PER06]. These examples suggest that the choice of operator is not merely an implementation detail but a fundamental modelling decision that encodes prior knowledge about the signal and the task.
Contributions
The main contributions of this paper are the following.
  • We introduce a principled taxonomy of five families of structured operators that extend or replace the convolution in learning-based image processing, covering decomposition-based, adaptive weighted, basis-adaptive, integral and kernel, and attention-based operators.
  • For each family, we provide a unified formal treatment, identifying the structural property of the convolution that each operator relaxes or replaces.
  • We present a comparative analysis of all families across multiple dimensions relevant to practitioners, including linearity, locality, equivariance, computational cost, and task suitability.
  • We discuss the open challenges in the area and highlight directions for future work, with particular attention to biomedical imaging and volumetric data.
Relation to own previous work
Several contributions in this taxonomy build upon and contextualise our own prior work. The decomposition-based family is grounded in our learning-based low-rank denoising approach [CP22], where a network predicts the optimal SVD thresholds for local image patches. The adaptive weighted family extends the framework of our weighted convolution with optimal density functions [CP25b, CP25c]. The basis-adaptive family encompasses our work on adaptive membership functions for the F-transform [CP24]. The computational aspects of operator optimisation are informed by our study of high-performance solvers for signal processing minimisation problems [CP25a]. Applications to medical imaging throughout the paper are motivated by our survey on ultrasound image processing [CP26].
2 Background: The Convolution Operator
We recall the definition and properties of the discrete convolution operator as it is used in CNNs, which serves as the common reference point throughout the taxonomy.
2.1 Definition
Given a 2D input signal $\bm{I}\in\mathbb{R}^{R\times C}$ and a kernel $\bm{w}\in\mathbb{R}^{K_{a}\times K_{b}}$, the discrete convolution is defined as
$(\bm{I}\ast\bm{w})_{ij}\;:=\;\sum_{a=1}^{K_{a}}\sum_{b=1}^{K_{b}}\bm{w}_{ab}\cdot\bm{I}_{i+a-\lfloor K_{a}/2\rfloor,\;j+b-\lfloor K_{b}/2\rfloor},$ (1)
for $i=1,\ldots,R$ and $j=1,\ldots,C$. In a CNN with $F$ output channels and $D$ input channels, a tensor of kernels $\bm{W}\in\mathbb{R}^{K_{a}\times K_{b}\times D\times F}$ is applied to produce the feature map $\bm{I}\ast\bm{W}$. The kernel weights are the trainable parameters of the layer, optimised by minimising a task-specific loss function via back-propagation [LBD+89, RUD16].
2.2 Structural Properties
The convolution operator has four structural properties that collectively define its expressive power and its limitations.
Linearity
The map $\bm{I}\mapsto\bm{I}\ast\bm{w}$ is linear in both $\bm{I}$ and $\bm{w}$. This enables efficient gradient computation but prevents the operator from modelling non-linear local interactions.
Translational equivariance
The same kernel $\bm{w}$ is applied at every spatial position, so that shifting the input produces a correspondingly shifted output: $T_{\delta}(\bm{I})\ast\bm{w}=T_{\delta}(\bm{I}\ast\bm{w})$, where $T_{\delta}$ denotes a spatial shift by $\delta$. This inductive bias is appropriate for natural images but may be suboptimal for structured data with position-dependent statistics, such as medical images with fixed anatomical priors.
Locality
The output at position $(i,j)$ depends only on the $K_{a}\times K_{b}$ neighbourhood $\mathcal{N}(\bm{I}_{ij})$. This limits the receptive field of each layer and requires deep stacking of layers to capture long-range dependencies.
Uniform weighting
The kernel applies the same learned weight to each relative position in the neighbourhood, irrespective of the local content of the signal. In particular, the kernel does not distinguish between positions that carry structural information (e.g., edges) and positions that carry noise.
2.3 Computational Cost
For an input of size $R\times C$, $F$ output channels, and kernel size $K\times K$, the computational cost of evaluating Eq. (1) is $\mathcal{O}(RCFK^{2})$. Practical implementations exploit FFT-based convolution [MHL13], the Winograd algorithm [LG16], and GEMM-based unrolling [Co14] to reduce the constant factor, while maintaining the same asymptotic complexity.
2.4 The Convolution as a Reference Point
Table 1 summarises the four structural properties discussed above. Each operator family introduced in Sections 37 relaxes or replaces one or more of these properties in a principled way, motivated by the structure of the signal and the requirements of the task. This framing provides the organising principle of the taxonomy.
Table 1: Structural properties of the standard convolution operator, used as a reference point for the taxonomy.
Property Convolution Implication
Linearity Yes Efficient gradient; no local non-linearity
Transl. equivariance Yes Position-independent feature extraction
Locality Yes (fixed $K$) Limited receptive field per layer
Uniform weighting Yes Content-agnostic neighbourhood averaging
3 Decomposition-Based Operators
Decomposition-based operators replace the convolution’s uniform averaging with a factorisation that explicitly separates the signal into structural components. The canonical example is the Singular Value Decomposition (SVD), which decomposes a matrix into orthogonal bases weighted by singular values, providing a natural separation between high-energy (structural) and low-energy (noise) components.
3.1 Local SVD Operators
Given a patch $\bm{P}\in\mathbb{R}^{m\times n}$ extracted from the neighbourhood of a pixel, the thin SVD is $\bm{P}=\bm{U}\bm{S}\bm{V}^{\top}$, where $\bm{U}\in\mathbb{R}^{m\times r}$, $\bm{S}=\mathrm{diag}(s_{1},\ldots,s_{r})$, $\bm{V}\in\mathbb{R}^{n\times r}$, and $r=\min(m,n)$. A denoised approximation is obtained by applying a threshold $\tau$ to the singular values:
$\hat{\bm{P}}=\bm{U}\,\mathrm{diag}(\tau(\bm{s}))\,\bm{V}^{\top},$ (2)
where $\tau(\bm{s})=\max(\bm{s}-\bm{\lambda},\bm{0})$ for a vector of thresholds $\bm{\lambda}$. This operation is non-linear (the threshold is a non-linear function of the singular values) and structure-aware (it explicitly separates the low-rank signal from the residual noise component).
The Weighted Nuclear Norm Minimisation (WNNM) [GZZ+14] applies this principle by combining SVD thresholding with block-matching to aggregate similar patches into a 3D stack, a strategy also used in BM3D [DFK+06]. Our learning-based extension [CP22] replaces the hand-crafted threshold selection with a network that predicts the optimal thresholds $\bm{\lambda}$ from the singular values of each block, trained to minimise the reconstruction error with respect to the ground-truth image. Algorithm 1 summarises this approach.
Algorithm 1 Learning-based SVD denoising [CP22]
1:Noisy image $\bm{Y}$, trained network $\mathcal{N}$, number of iterations $T$
2:Denoised image $\hat{\bm{X}}$
[…]
4:for $t=1,\ldots,T$ do
5:  Extract 3D blocks $\{\bm{B}_{k}\}$ via block matching on $\hat{\bm{X}}^{(t-1)}$
6:  for each block $\bm{B}_{k}$ do
7:   Compute $\bm{U}_{k},\bm{S}_{k},\bm{V}_{k}\leftarrow\mathrm{SVD}(\bm{B}_{k})$
8:   Predict thresholds $\hat{\bm{\lambda}}_{k}\leftarrow\mathcal{N}(\mathrm{diag}(\bm{S}_{k}))$
9:   $\hat{\bm{B}}_{k}\leftarrow\bm{U}_{k}\,\mathrm{diag}(\bm{S}_{k}-\hat{\bm{\lambda}}_{k})\,\bm{V}_{k}^{\top}$
10:  end for
11:  $\hat{\bm{X}}^{(t)}\leftarrow\mathrm{Aggregate}(\{\hat{\bm{B}}_{k}\})$
12:end for
13:return $\hat{\bm{X}}^{(T)}$
3.2 Tensor Decomposition Operators
For volumetric data and multi-channel images, the Higher-Order SVD (HOSVD) [KB09] generalises the matrix SVD to tensors $\mathcal{X}\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ via the Tucker decomposition $\mathcal{X}\approx\mathcal{G}\times_{1}\bm{U}^{(1)}\times_{2}\bm{U}^{(2)}\cdots\times_{d}\bm{U}^{(d)}$, where $\mathcal{G}$ is the core tensor and $\bm{U}^{(k)}$ are mode-$k$ factor matrices. Applied locally to image patches, HOSVD simultaneously exploits spatial, channel, and depth correlations, making it particularly effective for hyperspectral and volumetric data [ZYL+19] .
3.3 Low-Rank Layer Approximations
Decomposition principles can also be applied at the layer level rather than the patch level. Low-rank approximations of the weight tensor [SKS+13, LGR+15] factorise $\bm{W}\approx\bm{W}_{1}\bm{W}_{2}$ to reduce the number of parameters and the computational cost. While this is primarily a compression technique, it encodes a structural prior (low-rank kernel) that can improve generalisation in data-scarce regimes.
3.4 Properties and Task Suitability
Decomposition-based operators break the uniform weighting property of the convolution by assigning different contributions to different spectral components. They are non-linear (due to thresholding) and content-adaptive (the decomposition depends on the local patch structure). They are most naturally suited to image-to-image tasks where noise removal, compression, or structured feature extraction is required [HLL18]. Their main limitation is computational cost: computing the SVD of each patch at every layer is significantly more expensive than a standard convolution.
4 Adaptive Weighted Operators
Adaptive weighted operators retain the local neighbourhood structure of the convolution but modulate the kernel weights as a function of spatial position, signal content, or an externally optimised density function. They relax the uniform weighting and, in some cases, the translational equivariance property.
4.1 Convolution with Density Functions
A density function $\bm{\Phi}\in\mathbb{R}^{K\times K}$ scales the kernel weights element-wise before applying the convolution. Given the standard kernel $\bm{w}$ and the density function $\bm{\Phi}$, the weighted convolution is defined as [CP25b]
$(\bm{I}\ast\bm{w}_{\bm{\Phi}})_{ij}\;:=\;\sum_{a,b}(\bm{\Phi}_{ab}\,\bm{w}_{ab})\,\bm{I}_{i+a-\lfloor K/2\rfloor,\,j+b-\lfloor K/2\rfloor}.$ (3)
The density function encodes a prior about the relative importance of pixels at different distances from the reference pixel. When $\bm{\Phi}=\bm{1}$, Eq. (3) reduces to the standard convolution.
The optimal density function is computed by solving an outer optimisation problem [CP25b]: given a learning model $\mathcal{M}_{\bm{\Phi}}$ that minimises the loss function $\mathcal{L}$ with respect to the kernel weights $\bm{W}$ for a fixed $\bm{\Phi}$, the optimal density function $\bm{\Phi}^{*}$ is found by minimising the value of this minimised loss over all admissible density functions:
$\bm{\Phi}^{*}=\arg\min_{\bm{\Phi}}\;\min_{\bm{W}}\mathcal{L}(\mathcal{T},\,\hat{\mathcal{T}}(\bm{W}_{\bm{\Phi}})).$ (4)
Crucially, this formulation separates the optimisation of the kernel weights (solved by stochastic gradient descent) from the optimisation of the density function (solved by a global derivative-free method such as DIRECT-L [CP25a, GK01]). Algorithm 2 outlines the procedure.
Algorithm 2 Optimal density function computation [CP25b]
1:Dataset $(\mathcal{I},\mathcal{T})$, kernel size $K$, bounds $\alpha_{\min},\alpha_{\max}]$
2:Optimal density function $\bm{\Phi}^{*}$
3:Initialise $\bm{\alpha}=[1,\ldots,1]$
4:repeat
5:  $\bm{\Phi}\leftarrow\bm{\alpha}\bm{\alpha}^{\top}$
6:  $\bm{W}^{*}\leftarrow\arg\min_{\bm{W}}\mathcal{L}(\mathcal{T},\hat{\mathcal{T}}(\bm{W}_{\bm{\Phi}}))$ (SGD)
7:  Evaluate $f(\bm{\alpha})\leftarrow\mathcal{L}(\mathcal{T},\hat{\mathcal{T}}(\bm{W}^{*}_{\bm{\Phi}}))$
8:  Update $\bm{\alpha}$ via DIRECT-L using $f(\bm{\alpha})$
9:until convergence
10:return $\bm{\Phi}^{*}\leftarrow\bm{\alpha}^{*}(\bm{\alpha}^{*})^{\top}$
4.2 Dynamic Convolution
Dynamic convolution [CDL+20b] aggregates multiple parallel kernels $\{\bm{w}^{(k)}\}_{k=1}^{K}$ with input-dependent attention weights $\pi_{k}(\bm{I})$:
$\bm{I}\ast\bm{w}_{\mathrm{dyn}}\;:=\;\bm{I}\ast\left(\sum_{k=1}^{K}\pi_{k}(\bm{I})\,\bm{w}^{(k)}\right).$ (5)
The attention weights $\pi_{k}$ are produced by a small auxiliary network conditioned on a global average-pooled representation of the input. This makes the effective kernel content-dependent while preserving the local structure and translational equivariance of the convolution. Omni-dimensional dynamic convolution [LZY22] extends this idea to multi-dimensional attention over spatial, channel, and filter dimensions simultaneously.
4.3 Deformable Convolution
Deformable convolution [DQX+17] learns a set of spatial offsets $\{\Delta p_{k}\}$ that shift the sampling locations of the kernel:
$(\bm{I}\ast_{\mathrm{def}}\bm{w})_{ij}\;:=\;\sum_{k}\bm{w}_{k}\cdot\bm{I}(p_{ij}+p_{k}+\Delta p_{k}),$ (6)
where bilinear interpolation is used for non-integer sampling positions. Deformable convolution relaxes both the uniform weighting and the locality property: the receptive field adapts to the geometric structure of the signal.
4.4 Properties and Task Suitability
Adaptive weighted operators are particularly effective in scenarios where the relative importance of neighbouring pixels varies across the image, such as tasks involving structured textures, edges, or anisotropic noise. Density function optimisation [CP25b, CP25c] improves convergence speed and final accuracy for both image-to-image (denoising, PSNR improvement of 6–7%) and image-to-label (classification, accuracy improvement of 7 percentage points) tasks, without increasing the number of trainable parameters. The computational overhead is modest ($\approx 7$% on GPU).
5 Basis-Adaptive Operators
Basis-adaptive operators define the analysis and synthesis bases as learnable or data-dependent objects, replacing the fixed Fourier-like bases implicit in the standard convolution.
5.1 F-Transform with Adaptive Membership Functions
The fuzzy transform (F-transform) [PER06] projects a signal $f$ onto a set of fuzzy partition functions $\{A_{k}\}_{k=1}^{n}$ defined over the signal domain. The $k$-th F-transform component is:
$F_{k}\;:=\;\frac{\sum_{x\in\Omega}f(x)\,A_{k}(x)}{\sum_{x\in\Omega}A_{k}(x)}.$ (7)
The standard F-transform uses fixed, regularly spaced membership functions. Our extension [CP24] optimises the membership functions $\{A_{k}\}$ jointly with the network weights, making the projection bases adaptive to the signal statistics. Algorithm 3 summarises the adaptive F-transform layer.
Algorithm 3 Adaptive F-transform layer [CP24]
1:Input signal $f$, initialised membership functions $\{A_{k}^{(0)}\}$, loss $\mathcal{L}$
2:Adapted membership functions $\{A_{k}^{*}\}$, output components $\{F_{k}\}$
3:repeat
4:  for each $k$ do
5:   $F_{k}\leftarrow\sum_{x}f(x)\,A_{k}(x)/\sum_{x}A_{k}(x)$
6:  end for
7:  Compute $\mathcal{L}$; backpropagate gradients w.r.t. $\{A_{k}\}$ and network weights
8:  Update $\{A_{k}\}$ and network weights via SGD
9:until convergence
10:return $\{A_{k}^{*}\}$, $\{F_{k}\}$
The F-transform provides a natural connection between fuzzy set theory and signal processing, allowing the incorporation of domain knowledge into the choice and shape of the partition functions. Unlike the convolution, which applies the same kernel at every position, the F-transform aggregates information over the entire domain with position-dependent weights, relaxing both the locality and the translational equivariance constraints.
5.2 Learnable Wavelet Transforms
Wavelets provide a multi-scale decomposition of the signal, capturing features at different resolutions. Learnable wavelet layers [LZL+19] parameterise the wavelet filters and optimise them end-to-end, combining the interpretability of wavelet theory with the flexibility of learned representations [RV17]. The shearlet transform [KLR16] extends wavelets with directional sensitivity, providing a richer basis for capturing anisotropic features such as edges and ridges. Applied to denoising, shearlets are particularly effective for images with strong directional structures.
5.3 Sparse Dictionary Learning
An over-complete dictionary $\bm{D}\in\mathbb{R}^{n\times M}$ ($M\gg n$) represents the signal as a sparse linear combination of atoms: $\bm{y}\approx\bm{D}\bm{a}$, $\|\bm{a}\|_{0}\leq s$. The K-SVD algorithm [AEB06] alternates between sparse coding (finding $\bm{a}$ given $\bm{D}$) and dictionary update (refining each atom via SVD of the representation error). Dictionary learning can be embedded into a CNN as a learnable layer, replacing the convolution with a pursuit-and-synthesis operation.
5.4 Properties and Task Suitability
Basis-adaptive operators relax the translational equivariance and uniform weighting properties. Their strength lies in tasks where the signal has a known or learnable structure in a transformed domain (e.g., sparsity in a wavelet basis, smoothness in a fuzzy partition). They are particularly useful in medical imaging, where physical acquisition models (e.g., ultrasound speckle, MRI $k$-space sampling) motivate specific choices of basis [CP26].
6 Integral and Kernel Operators
Integral and kernel operators generalise the convolution by allowing the kernel to depend on the absolute or relative position of the pixels, rather than only on their relative offset. This relaxes the translational equivariance property.
6.1 Non-Local Means and Non-Local Neural Networks
The Non-Local Means (NLM) filter [BCM05] computes the output at position $i$ as a weighted average over all positions $j$:
$\hat{f}(i)=\frac{1}{Z(i)}\sum_{j}w(i,j)\,f(j),\quad w(i,j)=\exp\!\left(-\frac{\|\bm{P}_{i}-\bm{P}_{j}\|^{2}}{h^{2}}\right),$ (8)
where $\bm{P}_{k}$ is the patch at position $k$ and $h$ is a bandwidth parameter. The weight $w(i,j)$ depends on the similarity between patches at $i$ and $j$, making the operator content-adaptive and non-local. Non-Local Neural Networks [WGG+18] embed this idea into a differentiable layer, computing pairwise affinities between all feature map positions.
6.2 Radial Basis Function Networks
Radial Basis Function (RBF) networks [BL88] define the output as a linear combination of radially symmetric basis functions:
$y(\bm{x})=\sum_{k=1}^{M}c_{k}\,\phi(\|\bm{x}-\bm{\mu}_{k}\|),$ (9)
where $\phi$ is a radially symmetric function (e.g., Gaussian, multiquadric), $\bm{\mu}_{k}$ are the centres, and $c_{k}$ are the coefficients. Adaptive kernel-based sampling [CP21] optimises the centres, widths, and coefficients of Gaussian kernels to approximate arbitrary signals, including 2D images and vector fields, providing a structured alternative to fixed-grid convolution. Applied to image super-resolution, RBF interpolation with data-dependent centres provides a structured alternative to pixel-shuffle upsampling, encoding smoothness priors that are not available to a standard transposed convolution [DLH+16].
6.3 Convolutional Kernel Networks
Convolutional Kernel Networks (CKNs) [MKH+14] replace the dot product in the convolution with a positive definite kernel function $k(\cdot,\cdot)$:
$(\bm{I}\ast_{k}\bm{w})_{ij}\;:=\;\sum_{a,b}k(\mathcal{N}(\bm{I}_{ij}),\,\bm{w}_{ab}),$ (10)
where the kernel $k$ is typically a Gaussian or polynomial function over patch space. CKNs provide a principled connection between deep learning and kernel methods [SS02], and can be extended to graph-structured signals [CLL+20a].
6.4 Position-Encoding Operators
Coordinate convolution (CoordConv) [LLM+18] augments the input feature map with explicit coordinate channels, allowing the network to break translational equivariance in a controlled way. This is particularly useful for tasks where the absolute position of features is informative, such as object detection and localisation.
6.5 Properties and Task Suitability
Integral and kernel operators are the most general family in the taxonomy. They can model arbitrary dependencies between input positions, at the cost of increased computational complexity (NLM requires $\mathcal{O}(N^{2})$ operations per pixel for a full non-local search window) and reduced inductive bias. They are most effective in tasks requiring long-range context [WGG+18] or tasks where the translational equivariance prior is inappropriate. A key challenge is efficient implementation on large inputs [CP25a].
7 Attention-Based Operators
Attention-based operators can be viewed as the extreme case of the integral operator family, where the kernel is entirely learned from the data and depends on the global content of the input. They relax all four structural properties of the convolution and have become the dominant operator in large-scale vision models [DBK+21, Vo17].
7.1 Self-Attention
The scaled dot-product self-attention mechanism [Vo17] computes:
$\mathrm{Attention}(\bm{Q},\bm{K},\bm{V})\;=\;\mathrm{softmax}\!\left(\frac{\bm{Q}\bm{K}^{\top}}{\sqrt{d_{k}}}\right)\bm{V},$ (11)
where the query $\bm{Q}$, key $\bm{K}$, and value $\bm{V}$ matrices are linear projections of the input features. Each output position attends to all input positions, with weights determined by the similarity between query and key vectors. Self-attention is not translational equivariant, not local, and not linear in the input (due to the softmax).
7.2 Spatial and Channel Attention
Spatial attention [WPL+18] learns a weight map over the spatial dimensions of the feature map, emphasising task-relevant regions. Channel attention [WWZ+20] learns weights over the channel dimension, selecting the most informative feature channels. Both can be applied as lightweight modules that augment a standard convolutional backbone without replacing it entirely.
7.3 Vision Transformers
Vision Transformers (ViT) [DBK+21] replace convolutional layers entirely with multi-head self-attention applied to non-overlapping patches (tokens). This architecture achieves state-of-the-art performance on large datasets but requires substantially more training data than CNNs due to its weaker inductive biases. NAFNet [CCZ+22] and similar architectures apply simplified attention mechanisms that reduce the computational cost of the softmax operation for image restoration tasks.
7.4 Attention vs. Structured Operators
While attention mechanisms achieve remarkable performance, their success comes with the cost of weak structural priors and high computational cost ($\mathcal{O}(N^{2})$ in sequence length). In contrast, the structured operators discussed in Sections 36 embed domain knowledge into the operator design, which can be beneficial in data-scarce regimes such as medical imaging, or when interpretability is required [CP26]. Importantly, attention and structured operators are not mutually exclusive: hybrid architectures that combine local structured operators with global attention modules represent a promising research direction.
8 Comparative Analysis
8.1 Structural Properties
Table 2 summarises the structural properties of all five operator families with respect to the four reference properties of the standard convolution.
Table 2: Comparison of operator families across structural properties and task suitability. I2I: image-to-image tasks (denoising, super-resolution). I2L: image-to-label tasks (classification, detection). $\circ$: partial; ✓: yes; $\times$: no.
Operator family Linear Transl. equiv. Local Unif. weight. Adapt. to content Comp. cost I2I I2L Key ref.
Standard convolution $\times$ $\mathcal{O}(K^{2})$ [LBD+89]
Decomp.-based (SVD) $\times$ $\times$ $\times$ $\mathcal{O}(K^{3})$ $\circ$ [GZZ+14, CP22]
Adapt. weighted (density) $\circ$ $\times$ $\circ$ $\mathcal{O}(K^{2})$ [CP25b]
Adapt. weighted (dynamic) $\times$ $\mathcal{O}(K^{2})$ [CDL+20b]
Basis-adaptive (F-transform) $\times$ $\times$ $\times$ $\mathcal{O}(n\cdot|\Omega|)$ $\circ$ [CP24, PER06]
Basis-adaptive (wavelet) $\times$ $\circ$ $\times$ $\circ$ $\mathcal{O}(N\log N)$ $\circ$ [KLR16]
Integral (NLM / non-local) $\times$ $\times$ $\times$ $\mathcal{O}(N^{2})$ $\circ$ [BCM05, WGG+18]
Integral (RBF / kernel) $\times$ $\times$ $\times$ $\mathcal{O}(M\cdot N)$ [MKH+14]
Attention (self-attention) $\times$ $\times$ $\times$ $\times$ $\mathcal{O}(N^{2})$ [Vo17]
8.2 Design Principle Perspective
Viewing the operators through the lens of the four structural properties reveals a clear design trade-off. Moving from the standard convolution toward attention-based operators, each property is progressively relaxed: uniform weighting first (adaptive weighted operators), then translational equivariance (basis-adaptive and integral operators), then locality (attention-based operators). This relaxation is accompanied by an increase in expressive power and a corresponding increase in computational cost and reduction in inductive bias.
The role of inductive bias
Strong inductive biases—such as locality, translational equivariance, and uniform weighting—reduce the number of parameters required to learn a good representation and improve generalisation in data-scarce regimes. Weak inductive biases, conversely, allow the model to learn arbitrary dependencies from large datasets. The appropriate choice of operator therefore depends on the availability of training data, the structural properties of the signal, and the requirements of the task.
Computational cost
As shown in Table 2, operators that relax the locality property (integral operators, attention) have a significantly higher computational cost than local operators. For practical deployment on large images or volumetric data, the efficiency of the operator is a critical consideration [CP25a]. Optimised GPU implementations and approximations (e.g., linear attention [KVP+20]) can reduce this cost, but at the expense of expressive power.
8.3 Task Suitability Summary
For image-to-image tasks, decomposition-based and basis-adaptive operators are particularly well-suited because they explicitly encode structural properties of natural images (low-rank structure, multi-scale sparsity) that are relevant to denoising and super-resolution. For image-to-label tasks, adaptive weighted and attention-based operators are more appropriate, as they can capture global contextual information that is relevant for recognition and classification [CP25c]. In both cases, the optimal choice of operator depends on the specific signal statistics, the available training data, and the computational budget.
9 Open Challenges and Future Directions
Combination of operators
Most existing work considers operators in isolation. A promising direction is the design of architectures that combine multiple operator families, exploiting their complementary strengths. For example, a decomposition-based layer at the input could remove noise before a standard convolutional backbone extracts features, or an attention module could complement local structured operators with global context.
Operator selection as a meta-learning problem
Given the diversity of operators and the task-dependence of their performance, a natural question is whether the choice of operator can itself be automated. Neural architecture search (NAS) [ZL16] techniques could be extended to search over operator families, not just architectural hyperparameters.
Extension to 3D and volumetric data
All operators discussed in this paper are presented in 2D for clarity, but most have natural 3D extensions. For volumetric medical data (CT, MRI, ultrasound), structured operators that exploit the anisotropy of the acquisition process (e.g., different in-plane vs. through-plane resolution) could provide significant advantages [CP26]. Preliminary work on 3D weighted convolution is available at [CP25b].
Interpretability and theoretical analysis
While the empirical performance of many operators has been demonstrated, their theoretical properties—approximation power, convergence of optimisation, generalisation bounds—are less well understood. Formal analysis of the optimisation landscape of the density function optimisation problem [CP25b] and the convergence of adaptive F-transform layers [CP24] are examples of the type of theoretical grounding that is needed across all operator families. Efficient solver design for the related minimisation problems also remains an open challenge [CP25a].
Biomedical and real-world applications
Medical imaging represents a particularly important application domain for structured operators, due to the structured noise models (speckle in ultrasound, Rician noise in MRI), the anisotropic acquisition geometry, and the frequent data scarcity [CP26]. The operator families discussed in this survey, particularly decomposition-based and basis-adaptive operators, are well aligned with these challenges and deserve further investigation in this context.
Hardware-aware operator design
The theoretical $50\%$ overhead of the weighted convolution is reduced to approximately $7\%$ on modern GPU hardware due to memory-level parallelism [CP25b]. More broadly, the computational properties of operators are increasingly architecture-dependent, motivating the design of operators that are co-optimised with the underlying hardware.
10 Conclusions
We have presented a systematic taxonomy of structured operators that extend or replace the standard convolution in learning-based image processing. The five families—decomposition-based, adaptive weighted, basis-adaptive, integral and kernel, and attention-based operators—differ in which structural properties of the convolution they relax, and are consequently suited to different signal types, task requirements, and computational constraints.
The central message of this survey is that the convolution, while highly effective in many settings, is not the only or always the optimal choice of operator for learning-based image processing. A principled selection of the operator family, guided by the structural properties of the signal and the requirements of the task, can lead to significant improvements in both accuracy and efficiency. We hope that this taxonomy provides a useful reference for researchers and practitioners in the field, and motivates further work on the design, analysis, and application of structured operators for deep learning.
Acknowledgements
SC is part of RAISE Innovation Ecosystem, funded by the European Union – NextGenerationEU and by the Ministry of University and Research (MUR), National Recovery and Resilience Plan, Mission 4, Component 2, Investment 1.5, project “RAISE – Robotics and AI for Socio-economic Empowerment” (ECS00000035).
Declarations
Competing interests The author declares no competing interests.
Availability of data and materials No datasets were generated or analysed during the current study.